OpenStudy (anonymous):

find the derivative of tan^-1(2x)

OpenStudy (anonymous):

\[\tan^{-1} (2x)\]

OpenStudy (anonymous):

I"m told its suposed to be\[\frac{ 2 }{ 4x ^{2}+1 }\]

OpenStudy (anonymous):

My problem is that i thought the formula for the inverse trig function was :

OpenStudy (anonymous):

\[\frac{ 1 }{ x ^{2}+1 }\]

OpenStudy (anonymous):

So by the chain rule wouldn't i take the derivative of tan and plug in 2x?

OpenStudy (anonymous):

making it \[\frac{ 1 }{ 2x ^{2}+1 }\] ?

OpenStudy (anonymous):

why is it 2 on the top and 4x^2 on the bottom?

OpenStudy (anonymous):

y=arc tan 2x then by the chain rule you can find easily

OpenStudy (turingtest):

\[\frac d{dx}\tan^{-1}u=\frac1{u^2+1}\cdot\frac{du}{dx}\]in your case, \(u=2x\)

OpenStudy (anonymous):

yeah so it would be :

OpenStudy (anonymous):

\[\frac{ 1 }{ 2x+1 } * 2 \]

OpenStudy (turingtest):

if \(u=2x\) then \(u^2=?\)

OpenStudy (anonymous):

which would give \[\frac{ 2 }{ 2x+1}\]

OpenStudy (anonymous):

i meant the x to be squared on the bottom there

OpenStudy (turingtest):

no, think it through...\[u=2x\implies u^2=(2x)^2=?\]

OpenStudy (anonymous):

Oh. i was just putting the squared on the x and not also the two. thanks!

OpenStudy (anonymous):

That makes much more sense now. Gave you a medal!

OpenStudy (turingtest):

thanks, happy to help :)

OpenStudy (anonymous):

if y=arc tan (2x) then tan y = 2x..

OpenStudy (anonymous):